OpenStudy (anonymous):

solve exactly. 10 = 2(1+1/(10^x)0 x= ?

OpenStudy (anonymous):

\[10 = 2(1+(1/10^x))\] \[4 = 1/10^x\] \[10^x = 1/4\] \[\log_{10}(10^x) = \log_{10}(1/4) \] \[x = \log_{10}(4) = -.6021\]

OpenStudy (anonymous):

I hope this makes sense. Feel free to ask me anything

OpenStudy (anonymous):

it does but wondering how you got to 4 = 1/10^x

OpenStudy (anonymous):

10 = 2+2/10^x

OpenStudy (anonymous):

8 = 2/10^x

OpenStudy (anonymous):

Oh, first, divide both sides by 2 so you're left with \[5 = 1+(1/10^x)\] Then subtracting one gives you \[4 = 1/10^x\]

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

that is where I messed up

OpenStudy (anonymous):

you are awesome thank you

OpenStudy (anonymous):

no prob